How do you solve for x and y in the equation x^2 - xy = 823?
Answer by mananth(12270)

x^2-xy=823
x(x-y)=823
the factors of 823 are 1 & 823 since 823 is a prime
suppose x=823
then 823(823-y)=823
823-y=1
y=822